A metal alloy weighing 3 mg and containing 40% iron is melted and mixed with 12 mg of pure iron. What percent of the resulting alloy is iron?
Answer by stanbon(75887)

A metal alloy weighing 3 mg and containing 40% iron is melted and mixed with 12 mg of pure iron. What percent of the resulting alloy is iron?
-------
Equation:
iron + iron = iron
0.40*3 + 1*12 = x(3+12)
-----
40*3 + 1200 = 100x(15)
120+1200 = 1500x
1320 = 1500x
x = 1320/1500
x = 0.88 = 88%
Cheers,
Stan H.
================